# Stratified Hazard Sampling: Minimal-Variance Event Scheduling for CTMC/DTMC Discrete Diffusion and Flow Models

### v1

@@ 3.4 Simulation of Inhomogeneous Poisson Processes and Markov Jump Processes
Simulating CTMCs and non-homogeneous Poisson processes (NHPPs) is a classical topic. Thinning is a standard exact technique for NHPP simulation (lewis1979simulation). Uniformization is widely used in Markov jump process inference and sampling, enabling algorithms that avoid expensive matrix exponentials while preserving correctness (rao2013mjp). These classical tools inform modern samplers for high-dimensional discrete diffusion and CTMC generative models, and motivate variance-reduction schemes tailored to discrete generative inference.
@@ 4 Stratified Hazard Sampling
(a) Time/step-based EM: at fixed steps $t_{k}$, jump decisions are made probabilistically from the current rate $\lambda(t_{k})$, producing irregular (sometimes clumpy) events and high variance.
(b) Stratified Hazard Sampling: accumulate cumulative hazard $S(t)=\int_{0}^{t}\lambda(s)\,ds$ and place unit-spaced boundaries with a single offset $\theta\!\sim\!U(0,1)$. A jump occurs deterministically when $S(t)$ crosses $\theta+k$. This preserves the expected jump count while reducing variance to the minimal bound $\leq 1/4$.
Figure 1: Standard step-based sampler vs. SHS (ours).
We propose Stratified Hazard Sampling (SHS), adapting stratified event simulation to the step-based samplers used in CTMC/DTMC discrete generative models. SHS replaces the independent per-step change decisions $B_{ik}\sim\mathrm{Bernoulli}(p_{ik})$ (Section 2) with a single-phase stratification in cumulative hazard/jump-mass space. Concretely, SHS maintains the cumulative mass $S_{i,k}$ from (3) and triggers a jump whenever $S_{i,k}$ crosses unit-spaced thresholds $\{\theta_{i}+m\}_{m\in\mathbb{Z}_{\geq 0}}$ for a single random phase $\theta_{i}\sim\mathrm{Uniform}(0,1)$.
@@
10:  $S_{i}\leftarrow 0$; $k_{i}\leftarrow 0$; Draw $\theta_{i}\sim\mathrm{Uniform}(0,1)$
@@
16:   if $S_{i}\geq\theta_{i}+k_{i}$ then {/*crossed the next hazard boundary*/}
@@
One boundary per step.
The complete pseudocode of SHS is shown in Algorithm 1. When $p_{ik}\leq 1$ (as in DTMC kernels and in CTMC $\tau$-leaping with sufficiently small $h$), at most one unit boundary can be crossed per step, so the if statement on Line 10 is sufficient. If an application allows $p_{ik}>1$, the update should use an extra while loop to handle multiple crossings.
Jump-count concentration.
$J_{i}\;=\;I_{i}+\mathbf{1}[\theta_{i}<f_{i}],$ (9)
so $J_{i}\in\{I_{i},I_{i}+1\}$ and the long-tail jump-count fluctuations of step-wise Bernoulli sampling are eliminated. As a result, SHS suppresses the characteristic under-edit (too few substitutions) and over-edit (cascading substitutions) failure modes that arise under uniform-noise initialization.
@@ 4.1 Theoretical Properties
SHS is built on a simple primitive: randomized rounding of a nonnegative real mass $S$ into an integer count $J\in\mathbb{Z}_{\geq 0}$ using a single uniform random variable. In SHS, this mass is the total cumulative hazard/jump mass $S_{i}^{\mathrm{tot}}$ at each position (Eq. (9)).
Proposition 4.1 (Unbiasedness).
Fix $S\geq 0$ and draw $\theta\sim\mathrm{Uniform}(0,1)$. Let $\{S\}=S-\lfloor S\rfloor$ denote the fractional part and define $J=\lfloor S\rfloor+\mathbf{1}[\theta<\{S\}].$ Then $\mathbb{E}[J]=S$.
Proposition 4.2 (Minimal variance).
With $J$ as in Proposition 4.1, we have
$\mathrm{Var}(J)=\{S\}\big(1-\{S\}\big)\ \leq\ \frac{1}{4}.$ (10)
@@
Write $S=I+f$ with $I=\lfloor S\rfloor$ and $f=\{S\}\in[0,1)$. Then $J=I+B$ where $B=\mathbf{1}[\theta<f]\sim\mathrm{Bernoulli}(f)$. Hence $\mathbb{E}[J]=I+f=S$ and $\mathrm{Var}(J)=f(1-f)\leq 1/4$. See Appendix B for the full proof (and Appendix A for Proposition 4.1).
@@
4.2 Lexically Constrained Safety via Phase-Allocated SHS
Algorithm 2 Safety-SHS with phase allocation (blacklist $\mathcal{B}$)
1:Input: ModelStep routine (as in Algorithm 1); blacklist $\mathcal{B}$; time grid $t_{k}=kh$; initial state $X\sim p_{0}$
2:Output: final sample $X$ at $t_{n}$ (i.e., $t=1$)
3:for $i=1$ to NN do
4:  $S_{i}\leftarrow 0$; $k_{i}\leftarrow 0$; Draw $\theta_{i}^{0}\sim\mathrm{Uniform}(0,1)$
5:end for
6:Compute severity scores and allocate phases at a reference step $k_{0}$ (default $k_{0}=0$).
7:for $i=1$ to NN do
8:  $(p_{ik_{0}},\,q_{k_{0},i})\leftarrow\textsc{ModelStep}(t_{k_{0}},X_{t_{k_{0}}},i)$
9:  $s_{i}(X)\leftarrow\mathbf{1}[X^{i}\in\mathcal{B}]+\sum_{v\in\mathcal{B}}q_{k_{0},i}(v\mid X)$
10:end for
11: Assign phases $\{\theta_{i}\}_{i=1}^{N}$ from $\{\theta_{i}^{0}\}_{i=1}^{N}$ according to Eq. (14) using $\{s_{i}(X)\}_{i=1}^{N}$.
12:for $k=0,1,\dots,n-1$ do
13:  for $i=1$ to NN do
14:   $(p_{ik},\,q_{k,i})\leftarrow\textsc{ModelStep}(t_{k},X_{t_{k}},i)$
15:   $S_{i}\leftarrow S_{i}+p_{ik}$
16:   if $S_{i}\geq\theta_{i}+k_{i}$ then
17:    Sample $X_{t_{k+1}}^{i}\sim q^{\textsc{mask}}_{k,i}(\cdot\mid X_{t_{k}})$
18:    $k_{i}\leftarrow k_{i}+1$
19:   else
20:    $X_{t_{k+1}}^{i}\leftarrow X_{t_{k}}^{i}$
21:   end if
22:  end for
23:end for
24:return $X_{t_{n}}$
We consider blacklist-based lexical constraints, where users specify a forbidden token set $\mathcal{B}\subset\mathcal{V}$. Let the safe event be
$\mathcal{A}=\{x\in\mathcal{V}^{N}:\forall i,\,x_{i}\notin\mathcal{B}\}.$ (11)
Our goal is to approximate the conditional terminal distribution $p_{1}(x\mid\mathcal{A})$. As discussed in Section 2.2, exact conditioning corresponds to a Doob $h$-transform, which in general modifies both (i) which transitions are taken and (ii) how frequently they are taken. In contrast, naive masking modifies only the destination distribution and keeps the base jump frequency unchanged, often producing “late masking” artifacts.
Risk score.
At time $t$ and state $x$, we define a per-position violation risk score $r_{i}(t,x)\in[0,1]$ as the forbidden mass under the (change-conditional) one-step proposal:
$r_{i}(t,x)=\sum_{v\in\mathcal{B}}q_{t,i}(v\mid x),$ (12)
where $q_{t,i}$ is the destination distribution conditional on changing the token (cf. (2) for DTMC kernels).
Phase timing without modifying base jump mass.
Let $\theta_{1}^{0},\dots,\theta_{N}^{0}\overset{\text{i.i.d.}}{\sim}\mathrm{Uniform}(0,1)$ be pre-sampled SHS phases. We construct a risk-to-phase assignment that allocates smaller phases to higher-risk positions, thereby triggering earlier edits in cumulative hazard/jump-mass space while keeping the original per-position masses $p_{ik}$ intact. That is, in the CTMC case, we keep $p_{ik}=h\lambda_{i}(t_{k},x)$, while in the DTMC case, we keep $p_{ik}=1-P_{k,i}(x_{i}\mid x)$.
Given a reference time $t_{0}$ (default $t_{0}=0$) and state $x=X_{t_{0}}$, define a violation severity score
$s_{i}(x)=\mathbf{1}[x_{i}\in\mathcal{B}]+r_{i}(t_{0},x),$ (13)
and let $\sigma$ be a permutation such that $s_{\sigma(1)}(x)\geq\cdots\geq s_{\sigma(N)}(x)$ (ties broken by index). Let $\theta_{(1)}^{0}\leq\cdots\leq\theta_{(N)}^{0}$ denote the ascending order statistics of $\{\theta_{i}^{0}\}_{i=1}^{N}$. We then assign
$\theta_{\sigma(j)}\leftarrow\theta_{(j)}^{0},\quad j=1,\dots,N.$ (14)
Thus, positions with larger violation severity receive smaller phase offsets, making their first SHS boundary (and hence first edit opportunity) occur earlier.
Masked destination (which transition).
On a jump event at position $i$, we sample the new token from the masked destination
$q^{\textsc{mask}}_{t,i}(v\mid x)\propto q_{t,i}(v\mid x)\,\mathbf{1}[v\notin\mathcal{B}].$ (15)
Safety-SHS sampler (phase-allocated).
We run the same SHS accumulation as Algorithm 1 with unchanged per-position jump mass $p_{ik}$, but we replace the i.i.d. phase initialization by the above phase allocation rule, and we sample destinations using $q^{\textsc{mask}}$. This is a drop-in modification that does not require retraining and introduces no new continuous hyperparameters.
Effect on “late masking”.
In SHS, the first jump at position $i$ occurs when the cumulative mass $S_{i,k}$ crosses $\theta_{i}$. By assigning smaller $\theta_{i}$ to high-risk positions, Safety-SHS resolves censorship-active positions earlier, reducing the failure mode where the rest of the sequence adapts to an unsafe token before it is removed. We validate this by measuring the time-to-first-safe-edit CDF and terminal distribution shift under censorship (Appendix D).
Remark (bias vs. conditioning).
Unlike the rate-reweighting view of the Doob transform, phase allocation does not explicitly modify the base masses $p_{ik}$; instead it biases the relative event scheduling across positions through the SHS phase offsets. Since our target under censorship is the conditional distribution $p_{1}(\cdot\mid\mathcal{A})$, some bias relative to the unconditional dynamics is expected and can be beneficial in mitigating late masking artifacts. Note that phase allocation breaks the i.i.d. Uniform phase initialization used in Propositions 4.14.2; accordingly, the unbiasedness guarantee is not intended to apply to the censored dynamics.
@@ 4.3 On Preserving Multi-modality: Decomposing Sampler Variance
$\text{Var}[Q(X_{1})]$ $=\underbrace{\mathbb{E}_{\mathcal{T}}[\text{Var}_{\mathcal{U}}(Q(X_{1})\mid\mathcal{T})]}_{\text{(Term 1: Model Expressiveness)}}$
$\quad+\underbrace{\text{Var}_{\mathcal{T}}[\mathbb{E}_{\mathcal{U}}(Q(X_{1})\mid\mathcal{T})]}_{\text{(Term 2: Sampler Instability)}}.$
@@
SHS contributes by minimizing Term 2: bounding $\text{Var}(J_{i})\leq 1/4$ ensures near-optimal paths ($J_{i}\approx I$ or $I+1$), driving $\text{Var}_{\mathcal{T}}[\mathbb{E}_{\mathcal{U}}(\cdot)]$ near zero. Thus, SHS enhances reliability without compromising the model’s expressive power.
@@ 5 Experiments
5.1 Setup
Our goal is to isolate the effect of event-time sampling on a continuous-time discrete generative model. Throughout this section, we keep the trained model (rate/velocity predictor) fixed and only change the sampling procedure: (i) the standard time-step Euler–Maruyama (EM) sampler, and (ii) our Stratified Hazard Sampling (SHS).
Few-step regime.
To stress-test trajectory instability, we vary the number of discretization steps $n\in\{256,64,16,8\}$ while keeping the same model and schedule.
5.2 Jump Timing Diagnostics in Hazard Space
$H(t)=\int_{0}^{t}\lambda(s)\,ds.$ (16)
We record the hazard-space jump locations $\{H_{k}\}_{k\geq 1}$, where $H_{k}$ is the cumulative hazard value at the $k$-th jump.
EM baseline: Erlang/Gamma-shaped locations.
Under the standard NHPP view, inter-event increments in hazard space satisfy $\Delta H_{k}\sim\mathrm{Exp}(1)$ i.i.d., so the $k$-th jump location becomes
$H_{k}=\sum_{i=1}^{k}\Delta H_{i}\sim\mathrm{Gamma}(k,1),$ (17)
which yields an Erlang-shaped distribution as $k$ increases. Empirically, we observe broad, heavy-tailed variability in early jump locations under EM sampling, consistent with the “under-/over-edit” instability in the noise-start setting.
SHS: stratified (bounded-support) jump locations.
Figure 2: Hazard-space jump locations. Histograms of $H_{k}$ (the cumulative hazard at the $k$-th jump). EM sampling exhibits Erlang/Gamma-shaped variability, while SHS produces stratified, bounded-support locations, demonstrating reduced trajectory randomness.
5.3 Quantized MNIST Reconstructions: Eliminating Under-Edit
We next visualize the practical effect of trajectory stabilization on a simple discrete generation task. Starting from a noise initialization, we run the sampler for $n\in\{256,64,16,8\}$ discretization steps and compare the resulting reconstructions.
Observation.
EM sampling often exhibits an under-edit failure mode in the few-step regime: some positions experience too few jumps, leaving noticeable $x_{0}$-derived noise even late in sampling. In contrast, SHS produces visibly cleaner samples at early stages and substantially reduces residual noise in low-$n$ runs. Qualitatively, SHS maintains legible digit structure even at $n=16$ and $n=8$, where EM outputs remain dominated by unresolved noise.
Figure 3: Quantized MNIST reconstructions under varying step budgets. EM vs. SHS at $n\in\{256,64,16,8\}$ discretization steps. SHS reduces the under-edit regime and yields cleaner early-stage reconstructions.
5.4 UDLM Text Generation under Few-NFE Budgets
We next scale the same sampler-only ablation to a lightweight language-generation setting based on the UDLM evaluation protocol. As in all experiments, we keep the trained continuous-time discrete model (rate/velocity predictor) fixed and only swap the inference procedure between (i) the standard time-step Euler–Maruyama (EM) sampler and (ii) Stratified Hazard Sampling (SHS) (schiff2025simple).
@@ 5.4 UDLM Text Generation under Few-NFE Budgets
For each compute budget $\mathrm{NFE}\in\{4,8,16,32,64\}$, we generate $64$ sentences per run and repeat the evaluation over $5$ random seeds. We report (i) the sample entropy (as logged by the UDLM evaluation), and (ii) generative perplexity (Gen. PPL) computed by scoring the generated sentences using a fixed GPT2-large evaluator (lower is better).
@@ Results.
Table 1 summarizes the mean and standard deviation over seeds, and Figure 4 plots Gen. PPL vs. NFE. SHS consistently improves Gen. PPL in the few-step regime, with the largest gains at very low NFE: at NFE=$4/8/16/32$, SHS reduces Gen. PPL by approximately $6.3\%/5.2\%/5.8\%/3.8\%$ relative to EM, respectively, while the gap narrows at NFE=$64$ (about $1.4\%$). This pattern is consistent with our claim that SHS primarily removes sampler-induced trajectory variance (irregular event timing / jump-count fluctuations), thereby mitigating under-edit behavior when the discretization budget is tight.
NFE EM SHS
@@
Figure 4: Gen. PPL vs. NFE (UDLM protocol). SHS improves generation quality most notably in the few-step regime, indicating reduced sampler-induced instability under tight compute budgets.

### v2

@@ 3.3 Simulation of Inhomogeneous Poisson Processes and Markov Jump Processes
Simulating CTMCs and non-homogeneous Poisson processes (NHPPs) is a classical topic. Thinning is a standard exact technique for NHPP simulation (Lewis and Shedler, 1979). Uniformization is widely used in Markov jump process inference and sampling, enabling algorithms that avoid expensive matrix exponentials while preserving correctness (Rao and Teh, 2013). These classical tools inform modern samplers for high-dimensional discrete diffusion and CTMC generative models, and motivate variance-reduction schemes tailored to discrete generative inference.
3.4 Constrained generation and token-level lexical filtering
Lexical constraints have a long history as decoding-time control mechanisms in sequence generation. For autoregressive models, lexically constrained decoding has been studied through exact or approximate search procedures such as Grid Beam Search (Hokamp and Liu, 2017) and Dynamic Beam Allocation (Post and Vilar, 2018), with later work improving practicality and throughput (Hu et al., 2019). In safety-oriented settings, a widely used baseline is hard filtering (banning) of disallowed tokens or phrases. However, empirical evidence suggests that simple banned-word strategies can be insufficient as a complete safety solution, highlighting the distributional mismatch introduced by naive filtering (Gehman et al., 2020).
Our work does not propose a new constraint satisfaction algorithm. Instead, we study how common destination-side filtering heuristics interact with step-based CTMC/DTMC samplers. In particular, when lexical filtering is applied but the underlying stay-vs.-replace masses (or escape rates) remain unchanged, sampling quality can become sensitive to under-/over-edit trajectory tails. SHS addresses this orthogonal instability by reducing sampler-induced variance in edit counts and timing, while remaining a drop-in replacement that can be composed with destination filtering (Section 2.2 and Section 5.2).
@@ 4 Stratified Hazard Sampling
(a) At fixed steps $t_{k}$, jump decisions are made probabilistically from the current rate $\lambda(t_{k})$, producing irregular (sometimes clumpy) events and high variance.
(b) Cumulative hazard $S(t)=\int_{0}^{t}\lambda(s)\,ds$ and unit-spaced boundaries placed with a single offset $\theta\!\sim\!U(0,1)$. A jump occurs deterministically when $S(t)$ crosses $\theta+k$.
Figure 1: Standard sampler vs. SHS (ours).
We propose Stratified Hazard Sampling (SHS), adapting stratified event simulation to the step-based samplers used in CTMC/DTMC discrete generative models. SHS replaces the independent per-step change decisions $B_{ik}\sim\mathrm{Bernoulli}(p_{ik})$ (Section 2) with single-phase stratification in cumulative hazard/jump-mass space. Concretely, SHS maintains the cumulative mass $S_{i,k}$ from equation 3 and triggers a jump whenever $S_{i,k}$ crosses unit-spaced thresholds $\{\theta_{i}+m\}_{m\in\mathbb{Z}_{\geq 0}}$ for a single random phase $\theta_{i}\sim\mathrm{Uniform}(0,1)$.
One boundary per step.
The complete pseudocode of SHS is shown in Algorithm 1. Since $p_{ik}\leq 1$ holds by definition for DTMC kernels and is a prerequisite for valid CTMC $\tau$-leaping, at most one unit boundary can be crossed per step, so the if statement on Line 16 is sufficient.
Composability with lexical filtering.
SHS modifies only the stay-vs-replace event scheduling and leaves the per-jump destination sampling unchanged. Therefore, it can be composed with common destination-side modifications such as blacklist filtering by simply replacing $q_{t,i}$ with the filtered destination $q^{(\rho)}_{t,i}$ in  equation 5 while keeping the same scheduler.
Jump-count concentration.
$J_{i}\;=\;I_{i}+\mathbf{1}[\theta_{i}<f_{i}],$ (6)
so $J_{i}\in\{I_{i},I_{i}+1\}$ conditional on the realized cumulative mass $S_{i}^{\mathrm{tot}}$, and the long-tail jump-count fluctuations of step-wise Bernoulli sampling are eliminated. Note that in the full model the masses $p_{ik}(x)$ are state-dependent, so $S_{i}^{\mathrm{tot}}$ itself is a random variable whose distribution may differ between SHS and standard sampling (see Appendix D.3 for discussion). As a result, SHS suppresses the characteristic under-edit (too few substitutions) and over-edit (cascading substitutions) failure modes that arise under uniform-noise initialization.
@@
10:  $S_{i}\leftarrow 0$; $m_{i}\leftarrow 0$; Draw $\theta_{i}\sim\mathrm{Uniform}(0,1)$
@@
16:   if $S_{i}\geq\theta_{i}+m_{i}$ then {/*crossed the next hazard boundary*/}
@@ 4.1 Theoretical Properties
SHS is built on a simple primitive: randomized rounding of a nonnegative real mass $S$ into an integer count $J\in\mathbb{Z}_{\geq 0}$ using a single uniform random variable. Writing $S=I+f$ with $I:=\lfloor S\rfloor$ and $f:=S-I\in[0,1)$ (a decomposition we use throughout), the rounding rule is $J=I+\mathbf{1}[\theta<f]$ for $\theta\sim\mathrm{Uniform}(0,1)$. In SHS, this mass is the total cumulative hazard/jump mass $S_{i}^{\mathrm{tot}}$ at each position (equation 6).
Proposition 1 (Unbiasedness).
Fix $S\geq 0$ and draw $\theta\sim\mathrm{Uniform}(0,1)$. Write $S=I+f$ with $I:=\lfloor S\rfloor$ and $f:=S-I\in[0,1)$. Define $J:=I+\mathbf{1}[\theta<f]$. Then $\mathbb{E}[J]=S$.
Proposition 2 (Minimal variance).
With $J$ as in Proposition 1 (i.e., $S=I+f$ with $f:=S-\lfloor S\rfloor$), we have
$\mathrm{Var}(J)=f(1-f)\leq\frac{1}{4}.$ (7)
@@
Write $S=I+f$ with $I=\lfloor S\rfloor$ and $f=S-I\in[0,1)$. Then $J=I+B$ where $B=\mathbf{1}[\theta<f]\sim\mathrm{Bernoulli}(f)$. Hence $\mathbb{E}[J]=I+f=S$ and $\mathrm{Var}(J)=f(1-f)\leq 1/4$. See Appendix C for the full proof (and Appendix B for Proposition 1).
@@
A tail event relevant to blacklist robustness.
Beyond variance reduction, SHS also optimally suppresses the “no-edit” tail event under a fixed realized cumulative mass: among integer-valued random variables with $\mathbb{E}[J]=S$, SHS attains the minimum possible $\mathbb{P}(J=0)$. This provides a useful lens for lexical filtering robustness, where performance can be dominated by under-edited positions. See Appendix E.3 (Proposition 4).
@@ 4.2 On Preserving Multi-modality: Decomposing Sampler Variance
$\text{Var}[Q(X_{1})]$ $=\underbrace{\mathbb{E}_{\mathcal{T}}[\text{Var}_{\mathcal{U}}(Q(X_{1})\mid\mathcal{T})]}_{\text{(Term 1: Model Expressiveness)}}+\underbrace{\text{Var}_{\mathcal{T}}[\mathbb{E}_{\mathcal{U}}(Q(X_{1})\mid\mathcal{T})]}_{\text{(Term 2: Sampler Instability)}}.$
@@
SHS contributes by minimizing Term 2: bounding $\text{Var}(J_{i})\leq 1/4$ conditional on $S_{i}^{\mathrm{tot}}$ ensures near-optimal paths ($J_{i}\approx I$ or $I+1$), substantially reducing $\text{Var}_{\mathcal{T}}[\mathbb{E}_{\mathcal{U}}(\cdot)]$. Thus, SHS enhances reliability without compromising the model’s expressive power.
Crucially, while strictly separating $\mathcal{T}$ and $\mathcal{U}$ is an approximation due to state-dependent rates, this decomposition clarifies our design intent: to remove sampler-induced noise (Term 2) while leaving the model’s expressive stochasticity (Term 1) intact.
@@ 5 Experiments
5.1 Text Generation
We evaluate SHS on two uniform-noise discrete diffusion language models—UDLM (Schiff et al., 2025) and GIDD (von Rütte et al., 2025a)—to test whether the variance-reduction benefit generalizes across model families. In both cases, we keep the trained model (rate/velocity predictor) fixed and only swap the inference procedure between the standard sampler and SHS.
@@ 5.1 Text Generation
For UDLM we use $\mathrm{NFE}\in\{4,8,16,32,64\}$; for GIDD, $\mathrm{NFE}\in\{16,32,64,128,256\}$. UDLM generates $1024$ sentences per run over $5$ random seeds for each configuration, while GIDD generates $1024$ sentences with a single seed. We report sample entropy and generative perplexity (Gen. PPL) scored by a fixed GPT2-large evaluator (lower is better). We additionally evaluate on the 3B-parameter uniform diffusion model of von Rütte et al. (2025b) with $\mathrm{NFE}\in\{4,8,16,32,64,128\}$, $1000$ sentences, and a single seed; Gen. PPL is scored by Qwen2.5-7B.
@@ Results.
Tables 13 and Figure 2 summarize the results. Across all models, SHS consistently improves Gen. PPL, with the largest gains in the few-step regime: on UDLM, SHS reduces Gen. PPL by $6.3\%$ at NFE==4, narrowing to $1.4\%$ at NFE==64; on GIDD, improvements range from $17.3\%$ (NFE==16) to $15.7\%$ (NFE==256); on the 3B model (von Rütte et al., 2025b), SHS reduces Gen. PPL by up to $7.4\%$, while the confidence-based adaptive sampler yields substantially higher Gen. PPL than either parallel sampler despite comparable position-update budgets. These gains persist across three independently trained models (110M–3B), confirming that the benefit stems from the sampler, not from model-specific artifacts.
NFE Standard SHS
@@
NFE Standard SHS
[…]
16 $189.95\pm 8.02$ $5.246\pm 0.043$ $0.223\pm 0.004$ $\bm{157.02\pm 11.05}$ $\bm{5.054\pm 0.069}$ $\bm{0.231\pm 0.005}$
32 $106.76\pm 4.79$ $4.670\pm 0.045$ $0.266\pm 0.004$ $\bm{89.49\pm 2.74}$ $\bm{4.494\pm 0.031}$ $\bm{0.272\pm 0.003}$
64 $73.15\pm 5.13$ $4.290\pm 0.071$ $0.292\pm 0.005$ $\bm{63.57\pm 2.54}$ $\bm{4.151\pm 0.040}$ $\bm{0.299\pm 0.004}$
128 $60.10\pm 2.88$ $4.095\pm 0.048$ $0.310\pm 0.005$ $\bm{52.23\pm 2.03}$ $\bm{3.955\pm 0.039}$ $\bm{0.313\pm 0.005}$
256 $59.26\pm 2.95$ $4.081\pm 0.050$ $0.310\pm 0.005$ $\bm{49.92\pm 1.53}$ $\bm{3.910\pm 0.031}$ $\bm{0.317\pm 0.004}$
[…]
NFE Standard Adaptive (von Rütte et al., 2025b) SHS
[…]
8 $5.823\pm 0.673$ $338.1\pm 58.5$ $7.164\pm 0.432$ $1292.0\pm 1.6$ $\bm{5.756\pm 0.613}$ $\bm{316.1\pm 1.1}$
32 $4.751\pm 0.686$ $115.7\pm 0.5$ $6.477\pm 0.813$ $649.8\pm 1.2$ $\bm{4.680\pm 0.695}$ $\bm{107.7\pm 0.2}$
128 $4.379\pm 0.696$ $79.8\pm 0.5$ $6.160\pm 1.249$ $473.3\pm 1.6$ $\bm{4.321\pm 0.797}$ $\bm{75.2\pm 2.3}$
Table 3: 3B uniform diffusion model (von Rütte et al., 2025b): text generation under NFE budgets. 1000 sentences, single seed. Gen. PPL is scored by Qwen2.5-7B (lower is better). Adaptive uses the confidence-based decoding of von Rütte et al. (2025b), updating $\lceil N_{\mathrm{seq}}/\mathrm{NFE}\rceil$ positions per step so that the total position-update budget is comparable to the parallel samplers.
(a) UDLM
(b) GIDD
(c) 3B uniform diffusion
Figure 2: Gen. PPL vs. NFE. SHS consistently improves generation quality across models and scales, with the largest gains in the few-step regime.
5.2 Robustness under Lexical Constraints
We next evaluate robustness under random vocabulary truncation, which provides a controlled lexical stress test. We follow the same UDLM evaluation and GPT2-large scoring pipeline as Section 5.1, but fix the discretization budget to $\mathrm{NFE}=64$ and vary the blacklist severity.
Protocol.
For each blacklist ratio $\rho\in\{0.0,0.1,0.3,0.5,0.7\}$, we sample a forbidden set $\mathcal{B}_{\rho}\subset\mathcal{V}$ uniformly at random with $|\mathcal{B}_{\rho}|=\lfloor\rho|\mathcal{V}|\rfloor$, and define the allowed vocabulary $\mathcal{V}_{\rho}=\mathcal{V}\setminus\mathcal{B}_{\rho}$. We initialize from the restricted uniform noise $X_{0}\sim\mathrm{Unif}(\mathcal{V}_{\rho})^{N}$ (Eq. equation 4). At each edit, we enforce the blacklist by renormalizing the destination distribution to $\mathcal{V}_{\rho}$ (Eq. equation 5), while keeping the model-predicted change masses unchanged. We compare the Standard step-based sampler and SHS under the same sampled blacklist $\mathcal{B}_{\rho}$ per run.
We generate 1024 sentences per run and repeat over 5 random seeds. As in Section 5.1, we report (i) the sample entropy logged by the UDLM evaluation and (ii) Gen. PPL computed by a fixed GPT2-large evaluator (lower is better).
Results.
Random vocabulary truncation degrades generation quality as $\rho$ increases, as expected due to a reduced feasible token set. However, SHS consistently yields lower Gen. PPL than Standard across all blacklist ratios, and the degradation with $\rho$ is more gradual under SHS (Table 4 and Figure 3). These robustness trends are consistent with SHS’s optimal suppression of the "no-edit" tail under a fixed cumulative mass (Appendix E, Proposition 4).
$\rho$ Standard SHS
$\downarrow Entropy Gen. PPL\downarrow$
0.0 $6.377\pm 0.026$ $183.6\pm 8.0$ $\bm{6.370\pm 0.035}$ $\bm{181.0\pm 7.3}$
0.1 $6.715\pm 0.008$ $193.7\pm 3.1$ $\bm{6.667\pm 0.009}$ $\bm{181.8\pm 3.7}$
0.3 $6.481\pm 0.007$ $231.4\pm 4.3$ $\bm{6.438\pm 0.018}$ $\bm{217.1\pm 1.7}$
0.5 $6.017\pm 0.018$ $207.8\pm 2.8$ $\bm{5.974\pm 0.017}$ $\bm{197.5\pm 1.3}$
0.7 $4.634\pm 0.030$ $261.8\pm 1.6$ $\bm{4.590\pm 0.026}$ $\bm{244.2\pm 4.4}$
[…]
Figure 3: Robustness under random vocabulary truncation on UDLM (Gen. PPL vs. blacklist ratio $\rho$; lower is better). SHS degrades more gradually as $\rho$ increases.